Evaluate the integrals.

Integral(x+t) dx

Grade:9

1 Answers

Harsh davda
10 years ago
if here t is constant we can write the above statement as =integral(x)dx +integral(t)dx... now applying the formula for integral(x^n)dx={x^(n+1)}/(n+1)+C.. ,the above statement becomes =(x^2)/2+C+tx+C` .... here as t was constant , it comes out of the integral and 1 is remained .. we can write it as x^0.. and now C+C` becomes another constant.. say `P` so your answer is... = (x^2)/2+tx+P.. thanx